OpenStudy (gabylovesyou):

Determine the factors of 3x2 + 10x − 8.

OpenStudy (goformit100):

@nader1 Come Bro

OpenStudy (anonymous):

did you see before?

OpenStudy (luigi0210):

You could use the quadratic formula, or graph it.

OpenStudy (anonymous):

all you need to do is multiply out the brackets of each of those, and see which gives you the answer from the question. Stop being lazy!! Remember: (A + B)(C + D) = AC + AD + BC + BD ie - multiply everything by everything -except- the term it shares a bracket with. Also don't forget the right 'sign' eg (3x − 8)(x + 1) = 3x*x + 3x*1 - 8*x - 8*1 = 3x^2 + 3x - 8x - 8 = 3x^2 - 5x - 8 See how adding (-8 * 1) is the same as subtracting (8*1)... or just subtracting 8.

OpenStudy (gabylovesyou):

yes is it (x + 4) (3x - 2)

OpenStudy (anonymous):

answer(3x − 2)(x + 4)

OpenStudy (gabylovesyou):

i was right! yaaay me

OpenStudy (anonymous):

the other one before is it correct did you see it?